OpenStudy (anonymous):

solve for y in the following question. x= (x+1)(x+y)

OpenStudy (anonymous):

Possible answers are A. y= -x2/x+1 B. y= x/x+1 C. y=x-1/x2 D. y=x-1/-x2

OpenStudy (anonymous):

Maybe A

OpenStudy (anonymous):

Could I have a medal. (if I got it right)

OpenStudy (anonymous):

It was correct. Thank u